A HORIZONTAL FORCE F=mg/3 IS APPLIED ON THE UPPER SURFACE OF A UNIFORM CUBE OF MASS m AND SIDE a WHICH IS RESTING ON A ROUGH HORIZONTAL SURFACE HAVING u=1/2.THE DISTANCE B/W LINES OF ACTION OF mg and normal reaction is?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear vardaan

let distance between line of action of mg and N is x

so verticle force balance 

N=mg

clearly mg/3 is less than the maximum static friction force .

so only mg/3 static friction force will act .

now torque balance from the center of the cube

mg/3 *a/2  + mg/3 *a/2  = N*x

mga/3 =N*x

mga/3 = mgx

 x =a/3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin